OpenStudy (anonymous):

-6x-5y=13 2x+5y=19 solve simultaneous equation using elimination method i need it with solution

OpenStudy (anonymous):

What would happen if you add the 2 equations together?

OpenStudy (anonymous):

it will be -4x=32 x=8

OpenStudy (anonymous):

Nearly -4x=32 x=?

OpenStudy (anonymous):

x=8

OpenStudy (anonymous):

No :( Check your answer. (-4)*8=?

OpenStudy (anonymous):

i think its x=-8 and y=7

OpenStudy (nirmalnema):

-6x-5y=13 2x+5y=19 adding both the equations you will get -4x=32 so x= -8 put in any of the equation you will get y=7

OpenStudy (nirmalnema):

yup @sheelo73mughal you are correct

OpenStudy (anonymous):

and if we do this equation by substitution method

OpenStudy (nirmalnema):

you will get the same ans.

OpenStudy (nirmalnema):

find out the value of y or x from both the equations and equate the actually its double substitution y=19-2x / 5 and y = -6x-13 / 5 from above equations from hbere also u will get x= -8 and y = 7

OpenStudy (anonymous):

You can practice linear equations on my site http://www.saab.org/mathdrills/lineq.cgi